# Leading Higher Derivative Corrections to Multipole Moments of Kerr-Newman Black Hole

### v2

@@ Acknowledgement
We are grateful to Iosif Bena, Pablo Cano, Pengju Hu and Daniel Mayerson for useful discussions. L. Ma and Y. Pang are supported by the National Key R&D Program No. 2022YFE0134300 and the National Natural Science Foundation of China (NSFC) Grant No. 12175164 and No. 12247103. L. Ma is also supported by Postdoctoral Fellowship Program of CPSF Grant No. GZC20241211 and the China Postdoctoral Science Foundation under Grant Number 2024M762338. H. Lü is supported in part by NSFC grants No. 11935009 and No. 12375052.
@@ Appendix C Equations of motion
In this section, we derive the field equations from the extended Einstein-Maxwell theory described by the Lagrangian by $\mathcal{L}(g_{{\mu\nu}},R_{\mu\nu\rho\sigma},\widetilde{F}_{\mu\nu},F_{\mu\nu})$ in $D=4$. Let us first define
@@
$0=E_{g,\mu\nu}$ ::=:= $\frac{1}{\sqrt{-g}}\frac{\delta{(\sqrt{-g}\cal L)}}{\delta g^{\mu\nu}}=P_{(\mu}^{\ \ \alpha\beta\gamma}R_{\nu)\alpha\beta\gamma}-\frac{1}{2}g_{\mu\nu}\mathcal{L}+2\nabla^{\sigma}\nabla^{\rho}P_{(\mu|\sigma|\nu)\rho}-\frac{1}{2}\widetilde{\mathcal{M}}_{(\mu}^{\ \ \alpha}\widetilde{F}_{\nu)\alpha}$ (86)
@@
$\Delta\mathcal{M}_{\mu\nu}$ $=$ $\sum_{i=1}^{8}c_{i}\mathcal{M}_{\mu\nu}^{(e,i)},\quad\widetilde{\mathcal{M}}_{\mu\nu}=\sum_{i=1}^{4}d_{i}\widetilde{\mathcal{M}}_{\mu\nu}^{(o,i)}\ ,$

### v3

@@ Acknowledgement
We are grateful to Iosif Bena, Pablo Cano, Pengju Hu and Daniel Mayerson for useful discussions. L. Ma and Y. Pang are supported by the National Key R&D Program No. 2022YFE0134300 and the National Natural Science Foundation of China (NSFC) Grant No. 12175164, No. 12247103 and No. 12447138. L. Ma is also supported by Postdoctoral Fellowship Program of CPSF Grant No. GZC20241211 and the China Postdoctoral Science Foundation under Grant No. 2024M762338. H. Lü is supported in part by NSFC grants No. 11935009 and No. 12375052.
@@ Appendix C Equations of motion
In this section, we derive the field equations from the extended Einstein-Maxwell theory described by the Lagrangian by $\mathcal{L}(g^{{\mu\nu}},R_{\mu\nu\rho\sigma},\widetilde{F}_{\mu\nu},F_{\mu\nu})$ in $D=4$. Let us first define
@@
$0=E_{g,\mu\nu}$ ::=:= $\frac{1}{\sqrt{-g}}\frac{\delta{(\sqrt{-g}\cal L)}}{\delta g^{\mu\nu}}=P_{(\mu}^{\ \ \alpha\beta\gamma}R_{\nu)\alpha\beta\gamma}-\frac{1}{2}g_{\mu\nu}\mathcal{L}+2\nabla^{\sigma}\nabla^{\rho}P_{(\mu|\sigma|\nu)\rho}-\frac{1}{2}\widetilde{\mathcal{M}}_{(\mu}^{\ \ \alpha}\widetilde{F}_{\nu)\alpha}$
@@
$\Delta\mathcal{M}_{\mu\nu}$ $=$ $\sum_{i=1}^{8}c_{i}\mathcal{M}_{\mu\nu}^{(e,i)},\quad\widetilde{\mathcal{M}}_{\mu\nu}=\sum_{i=1}^{4}d_{i}\widetilde{\mathcal{M}}_{\mu\nu}^{(o,i)}\ ,$ (70)